[Hallc_running] [New Logentry] RC report Nov-28

berdnik at jlab.org berdnik at jlab.org
Sun Nov 28 13:30:03 EST 2021

Logentry Text:

We are done with data taking at the Q2=6.0, W=2.40, TH_shms=11.12deg (center) setting. We have  pi- 12932 events. https://logbooks.jlab.org/entry/3952018

We lost HMS Q3 for the second part of the day shift and for first quarter of the swing shift yesterday. Joe reset interlock in the hall by power cycling the magnet power supply during escorted control access. The issue cost us ~ 5h 37min of downtime.  https://logbooks.jlab.org/entry/3951535 and  https://logbooks.jlab.org/entry/3951658  and https://logbooks.jlab.org/entry/3951659

After we returned to physics production on LD2 till 12-30 today. 
We wasn`t able to rotate SHMS remotely and had a short escorted controlled access and Steve reset rotation interlock by power cycling the power supply.  The issue cost us ~ 59 min of downtime.   https://logbooks.jlab.org/entry/3951982

SHMS rotated to 13.12 degree https://logbooks.jlab.org/entry/3952014
We started data taking for Q2=6.0, W=2.40, TH_shms=13.12deg (left) setting... 

Run plan:

-- 13.12deg: 57hrs from Sunday at 13:30 is Tuesday at 19:30, plus 3hrs of Dummy goes to Wed at 01:30.

Pending issues:


RC meeting:

NO meeting today (Sunday)


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3952026
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211128/c8333360/attachment.html>

More information about the Hallc_running mailing list